### v1

# Afford2Act: Affordance-Guided Automatic Keypoint Selection for Generalizable and Lightweight Robotic Manipulation

Abstract
Vision-based robot learning often relies on dense image or point-cloud inputs, which are computationally heavy and entangle irrelevant background features. Existing keypoint-based approaches can focus on manipulation-centric features and be lightweight, but either depend on manual heuristics or task-coupled selection, limiting scalability and semantic understanding. To address this, we propose Afford2Act, an affordance-guided framework that distills a minimal set of semantic 2D keypoints from a text prompt and a single image. Afford2Act follows a three-stage pipeline: affordance filtering, category-level keypoint construction, and transformer-based policy learning with embedded gating to reason about the most relevant keypoints, yielding a compact 38-dimensional state policy that can be trained in 15 minutes, which performs well in real-time without proprioception or dense representations. Across diverse real-world manipulation tasks, Afford2Act consistently improves data efficiency, achieving an 82% success rate on unseen objects, novel categories, backgrounds, and distractors. Visualizations available at: https://afford2act.github.io/
I Introduction
Learning effective robot manipulation policies hinges on overcoming a fundamental representation challenge: how to extract the manipulation-related features which are both compact and expressive? Dense visual inputs such as images [1], pointclouds [2], or semantic dense features [3, 4] provide rich details, but burden a policy with extraneous information. In contrast, 2D keypoint [5] offers a sparse state representation that highlights only the essential features of objects’ motions while still being representative enough to learn a good policy. This sparsity directly enables a lightweight policy with far fewer but important features than dense representation and support effective and real-time inference with less computational demand. However, the real challenge is in discovering these keypoints automatically and reliably: picking too few or irrelevant points can omit critical information, while relying on human insight for what points to use does not scale. In this paper, we present a method for identifying task-relevant object points that are both minimal (compact) and sufficient (expressive) for effective robot manipulation.
The problem of identifying keypoints has received a lot of attention recently. Some previous work utilizes manual selection [6] and prompts large Vision-Language Models (VLMs) to mark affordances on images [7, 3]. While these methods can provide useful cues, they depend on human guidance or external knowledge, making them labor-intensive and domain-specific. At the other end of the spectrum, some approaches forgo sparsity altogether in favor of dense feature grid-based representations. For instance, all previous work [8, 9, 10, 11] learn from dense grid of points in the scene or the object. While dense features are expressive, they are not compact and incur heavy computational demands. Furthermore, dense features may also entangle irrelevant dynamics (such as background features), which limits scalability and focus.
Figure 1: Top: policy trained with original prompts (“hold”, “stir”) executes the actions over time ($t$). Middle: at test time, the same policy generalizes to synonym prompts (“pick”, “mix”). Bottom: affordance-guided keypoints with learned importance (green; variable count; ring radius $\propto$ attention) remain stable under unseen object shapes, lighting conditions, dynamic distractors, and scene clutter—leading to successful rollouts.
Figure 2: Overview of the Afford2Act pipeline. We extract task-relevant object keypoints using affordance masks and DINO filtering, track them across frames, and embed them with a transformer and attention gating. The policy head predicts pose and gripper commands from these compact keypoint representations, enabling vision-based manipulation without requiring proprioception.
Recently, there has been work on automatically selecting a subset of keypoints from a dense grid [12, 13, 14] in an end-to-end manner. In particular, ATK [12] jointly optimizes a masking model for selecting a compact keypoint set and a task-specific control policy, yielding strong performance on challenging visuomotor tasks. However, there are two key limitations in ATK that prohibit it from generalizing: (i) Joint training couples the masking model to the task, so extending to a new task requires retraining both components (e.g., moving from mug picking from left to right); (ii) Existing evaluations emphasize robustness to positional variation and visual perturbations, whereas other generalization regimes, duplicate objects and cross-category transfer (e.g., from mugs to teapots), have received limited attention. In summary, prior works either require significant prior knowledge (manual keypoints), incur high computational overhead (dense grids), or struggle with stability and semantic generalization. This state of affairs motivates a core question for robot learning: How can a robot learn to identify the most relevant interaction points, the task hotspots, directly from visual inputs?
To address this challenge, we propose Afford2Act: Affordance-Guided Keypoints to Act, which uses text-prompted affordance cues to automatically select task-relevant keypoints, with no manual labels. This pipeline contains three stages: (i) Affordance Filtering (e.g., “hold” $\rightarrow$ mug handle), (ii) Keypoint Pool Construction from a small set of masked DINO anchors, and (iii) Keypoint-Based Policy Learning, where a transformer-based encoder followed by a gating mechanism learns the importance of each keypoint. Because mask support and geometry vary, the number of selected keypoints per object also varies; thus, correspondence is defined as a flexible set of affordance-region points. As visualized in Fig. 1, the policy typically attends to a small effective subset (Fig. 11) of keypoints. With a 38-dimensional input, policies remain robust under open-vocabulary prompts and scene changes across six real-world tasks. Evaluating across six real-world tasks, Afford2Act consistently outperforms keypoint-selection baselines and RGB/RGB-D policies, and stays robust and generalizable to synonym prompts, lighting, distractors, and clutter (Fig. 1).
Contributions. We present:
  1. 1.
    A unified Affordance-Guided Keypoint Distillation framework that extracts sparse, semantically meaningful 2D keypoints directly from text prompts.
  2. 2.
    An efficient, attention-based framework that trains a compact keypoint-guided policy in $\sim$15 minutes.
  3. 3.
    Extensive real-robot experiments in six tasks, demonstrating superior 82% success in unseen instances and robustness to open-vocabulary prompts (Fig. 1).
Figure 3: Real-world rollouts of Afford2Act on Pour (top) and Cut (bottom) tasks. In each sequence (left to right: approach $\rightarrow$ interaction $\rightarrow$ completion), green dots mark keypoints automatically selected on the affordance region (e.g., knife for “cut”, handle/spout for “pour”). The translucent ring size reflects the attention weight at each step. While keypoints remain fixed on task-relevant parts, the policy dynamically shifts its focus: for “cut”, the attention moves from the blade tip toward the handle; for “pour”, from the edge to the main handle. This adaptive attention enables robust, context-aware manipulation.
II Related Work
Visual Representations for Policy Learning: Choosing the right visual representation is central to building robust manipulation policies. Classic approaches rely on raw sensory data which includes images, depth maps, or point clouds to provide a complete view of the environment [1, 2, 15]. While these representations are rich, they also saddle policies with high-dimensional inputs, requiring heavy data regimes and often resulting in policies that overfit to appearance rather than task. As a remedy, manipulation-centric and object-centric representations [3, 16, 17, 18, 19, 20, 21] filter out much of the visual clutter, focusing policy learning on the handful of features that matter most for action. However, these strategies often rely on hand-tuned detectors or dense input from images 3D reconstructions, limiting scalability and flexibility. Our method bypasses these constraints, extracting a compact, affordance-driven set of 2D keypoints that serve as a focused visual state, requiring neither dense geometry nor extensive supervision.
Keypoints and Automatic Anchor Discovery: Keypoints, when properly chosen, offer a natural low-dimensional representation that encodes an object’s essential affordances [6, 10, 11, 14]. Earlier work often assumed keypoints could be specified by experts, or required laborious labeling to maintain semantic consistency across objects [6, 22]. More recent efforts automate this selection, using object detector [10, 11] or learning which points are most predictive for the task at hand [12, 8]. For keypoint sampling from detected objects, the representations are still dense and lack semantic meaning. ATK, for instance, integrates keypoint selection with policy optimization, but can produce unstable and semantically ambiguous anchors. VLM-driven pipelines like MOKA [7] bring zero-shot power to keypoint discovery, but still rely on iterative visual prompting or domain knowledge. Our approach sidesteps both manual bias and unstable end-to-end discovery by using affordance cues as a reliable guide, automatically surfacing keypoints that are functionally meaningful and consistent, without relying on heavy annotation or black-box prompts.
Affordance Reasoning in Manipulation: The concept of affordances has become foundational in robotics, bridging perception and action by focusing on what an object “invites” a robot to do [23, 24]. Early works constructed dense affordance maps [25, 26, 27] or trained region classifiers, but these methods faced challenges in scaling and generalization. The rise of vision-language models and foundation models [7, 28] brought a new wave of affordance reasoning, inferring actionable regions from text and image cues. However, the leap from model outputs to actionable robot inputs is nontrivial: such approaches can be coarse, ambiguous, or computationally heavy. In contrast, we use affordance signals to guide the distillation of sparse, actionable keypoints, ensuring the resulting representation is both semantically grounded and directly useful for policy learning, combining the semantic strengths of foundation models with the simplicity and efficiency of classic keypoint methods.
III Afford2Act
Overview. Given single-view demonstrations $\mathcal{D}=\{(I^{(d)}_{1:T_{d}},\,\mathbf{u}^{(d)}_{1:T_{d}},\,A^{(d)})\}_{d=1}^{D}$, where $I^{(d)}_{t}\in\mathbb{R}^{H\times W\times 3}$ are images, $\mathbf{u}^{(d)}_{t}\in\mathbb{R}^{m}$ are actions (e.g., 6-DoF deltas and a gripper scalar), and $A^{(d)}$ is a text prompt, Afford2Act: (i) localizes the actionable region via a prompt-conditioned affordance mask (Section III-A), (ii) selects a small set of semantic keypoints on the reference frame (Section III-A), (iii) matches them within the mask on new instances to establish flexible correspondence ($\leq 15$) for various geometry and lightening (Section III-A), (iv) jointly tracks selected keypoints and encodes them into a transformer with a gating layer that learns phase-adaptive weights to encode a compact scene embedding for policy learning (Section III-B). The overall framework is shown in Fig. 2.
Formal Setup. Each demonstration $d$ is summarized by trajectories of $K$ 2D keypoints $\mathbf{P}^{(d)}=\{(x^{(d)}_{t,i},\,y^{(d)}_{t,i})\}_{t=1:T_{d},\;i=1:K}$, where indices $i$ refer to functional parts within the affordance region (e.g., “handle tip,” “blade edge”). This sparse, semantic representation lets the policy reason over roles rather than raw pixels.
III-A Affordance-Guided Keypoint Representation
Our pipeline for extracting these keypoints proceeds in four stages, all fully automatic and requiring no manual annotation (See Fig. 2).
Affordance Localization: Given the first frame $I_{1}$ of a demonstration affordance prompt, we use One-Shot Open Affordance model  [28] $f_{\text{aff}}$ to produce a heatmap: $H=f_{\text{aff}}(I_{1},A)\in[0,1]^{H\times W}$. This heatmap indicates regions of the scene that likely correspond to the intended action. To focus only on the actionable surface, we apply a robust quantile threshold (along with mild smoothing) to obtain a binary mask: $M(x,y)=\mathbb{1}\big[H(x,y)\geq\tau_{q}(H)\big]$. This mask captures fine structures such as a knife blade or cup rim and robustly filters out clutter. Text-based prompts provide a strong semantic prior, allowing the same model to adapt affordance detection across novel tools or object categories with minimal (such as the examples in Fig. 3).
Reference Keypoint Extraction: On a reference frame (in our case, the first frame of the first demonstration), $I_{1}^{r}$, we compute dense feature maps using a Vision Transformer backbone (ViT/DINO) [29]: $F_{r}=\phi(I^{r}_{1})\in\mathbb{R}^{H\times W\times D}$. We then restrict attention to the ROI defined by the affordance mask $M$ and cluster the features within that region to obtain $N$ reference anchor points:
$\mathcal{A}_{r}=\left\{(p_{i}^{r},\,\mathbf{f}_{i}^{r})\right\}_{i=1}^{N},\qquad p_{i}^{r}\in\Omega(M),\ \mathbf{f}_{i}^{r}=F_{r}(p_{i}^{r})$
This step is fully automatic and scales to new categories, in contrast to manual annotation-based approaches [6, 7].
Cross-instance Correspondence: Given a new demonstration, we extract features from its first frame, $F_{t}=\phi(I^{t}_{1})$. From each keypoint, $(p_{i}^{r},\mathbf{f}_{i}^{r})$ in the reference anchor, $\mathcal{A}_{r}$, we compute a cosine similarity map restricted to the affordance ROI in the new frame:
$S_{i}(x,y)=\left\langle\frac{\mathbf{f}_{i}^{r}}{\|\mathbf{f}_{i}^{r}\|},\,\frac{F_{t}(x,y)}{\|F_{t}(x,y)\|}\right\rangle,\quad(x,y)\in\Omega(M_{t})$
We assign the keypoint with the highest similarity,
$p^{t}_{i}=\arg\max_{(x,y)\in\Omega(M_{t})}S_{i}(x,y),\quad c^{t}_{i}=\max_{(x,y)\in\Omega(M_{t})}S_{i}(x,y)$
This matching ensures that the index $i$ refers to the same functional part (e.g., “handle tip”) across all demonstrations, even as object geometry and appearance vary. By constraining matches to the affordance mask, we avoid drifting to irrelevant but visually similar regions.
Temporal Keypoint Tracking: Once all $K$ keypoints are initialized on the first frame, we track them across the demonstration using a transformer-based tracker such as CoTracker [30]: $\{p_{t,i}\}_{t=1:T,\,i=1:K}=g_{\text{trk}}(I_{1:T},\,\{p^{1}_{i}\}_{i=1}^{K})$
Joint tracking improves robustness to occlusion and viewpoint changes, leveraging correlations among keypoints. The output is a set of trajectories in a semantic space where each index $i$ always refers to the same actionable region.
By the end of this process, every demonstration is described by a trajectory of $K$ task-relevant, affordance-grounded keypoints. This sparse and semantically aligned representation serves as the input to our downstream policy (Section III-B), bridging the gap between the efficiency of keypoints and the semantic expressiveness of affordances, without requiring heavy 3D pipelines or manual keypoint curation.
III-B Policy Design
At each interfere step $t$, taking the $K$ keypoints from keypoint distillation $\{p_{t,i}\}_{i=1}^{K}$, the question becomes: how do we architect a policy, $\pi_{\theta}$, that prioritizes the most task-relevant keypoints? After embedding each keypoint, $z^{1}=[\phi_{\theta}(p_{1})\dots\phi_{\theta}(p_{K})]^{\top}$, we then compute contextualized token features using a Transformer chunk, $z=\mathcal{T}_{\theta}(z_{1})$, similar to [31, 32] so that each keypoint is interpreted in the context of the others (e.g., a keypoint on the blade of a knife carries a different affordance than one on the handle). Since the number of keypoints $m$ can exceed the task-relevant cardinality, we introduce a gating network, $g_{\psi}$, that predicts a non-negative importance weight for each keypoint before pooling them into a single scene feature vector $h$:
$\hat{w}=g_{\psi}(z)=\operatorname{softmax}(w_{g}^{\top}z+b_{g}),\hat{z}=\hat{w}^{\top}z,\quad h=\frac{1}{K}\sum_{i=1}^{K}\hat{z}_{i}$
Then, the scene embedding, $h$ is passed through a shared action head that is later split into two linear layers: one for the delta actions and the other for the gripper state. While we use a two-layer MLP, the head is modular and can be replaced by a generative policy head [33, 34, 1].
IV Experiments
Our experiments are designed to evaluate three claims:
  1. 1.
    Keypoint Effectiveness: Can a policy trained exclusively on a compact set of 2D keypoints outperform policies trained on full images, dense features, and even a large vision-language foundation model?
  2. 2.
    Keypoint Priors: Which alternative keypoint selection strategies (e.g., using object bounding box, uniform grids of points, or manually selected points) can match the informativeness of our affordance-guided keypoints?
  3. 3.
    Generalization: How well does our policy generalize to new scenarios, including novel objects within seen categories, entirely new object categories, the presence of distractor objects, and other scene variations?
Additionally, we perform targeted ablation studies to test the robustness to different language prompts, to verify the necessity of the gating network, and to explore an extended two-stage pipeline for long-horizon tasks.
Figure 4: (Left) A figure of our physical setup, composing of a UR3e Robotic arm and one side RealSenseD435i Camera. (Right) shows the objects we used for seen (right) and unseen (left) scenarios.
Figure 5: Qualitative generalization on two tasks. Each row shows rollouts for Pour (top) and Kick (bottom). Left: training objects (seen). Right: held-out settings with unseen scenes (new backgrounds/tabletops) and, for Pour, unseen object instances. Within each panel we show three key frames (setup, approach, completion). Without any test-time adaptation, Afford2Act trained from single-view RGB using affordance-guided keypoints successfully transfers by attending to the functional part (mug handle / ball) while ignoring clutter.
IV-A Setup and Tasks
Our method is driven by task semantics expressed as natural-language prompts that condition an affordance model. We focus on these core affordances—hold, cut_with, brush_with, pour, and kick—and systematically construct a variety of manipulation tasks by pairing each affordance with diverse objects and contexts. The mapping from affordances to example tasks is as follows:
Hold: Grasp and lift an object with a handle (e.g., a mug) using the hold affordance. (Fig. 1)
Cut: Perform a cutting motion on an object using a knife, guided by the cut_with affordance. (Fig. 3)
Brush: Sweep debris using a hand brush. This task leverages both the hold and brush_with affordances in coordination (holding the brush and using it to brush). (Fig. 2)
Pour: Execute a pouring motion with a vessel (e.g., tipping a cup or pitcher) using the pour affordance. (Fig. 5)
Stir: Stir the contents of a bowl using a spoon or similar tool, guided by a combination of hold (to grasp the spoon) and a stir affordance. (Fig. 1)
Kick: Push a ball into a goal using the kick affordance. 5
For each task, we train a policy and test it with a variety of seen and novel objects (see Fig. 4). The objects are physically diverse in size, shape, appearance, and material. This diversity allows us to evaluate generalization at both the instance level (new objects of a trained category) and the category level (objects from entirely new categories). We collected 40 teleoperated demonstrations per task for training, with object positions randomized across demonstrations.
IV-B Input Modality Baselines
We benchmark our approach against four baseline methods to evaluate the keypoints effectiveness:
  • RGB-BC: A behavior cloning policy trained on raw RGB images, using a ResNet-18 visual encoder and a multilayer perceptron (MLP) for the action output.
  • RGB-D-BC: The framework as RGB-BC, but trained on 4-channel RGB-D images (RGB plus monocular depth from an Intel RealSense camera).
  • UAD-2D-BC: The same framework as RGB-BC, but the images are replaced with the affordance mask generated by UAD [3]. This baseline can verify the effectiveness of sparse keypoints rather than dense representations.
  • Foundation Model (Pi0): A large vision-language-action foundation model (Pi0) [35] fine-tuned on our demonstration data. We evaluate this model with a single side camera on a UR3e robot arm, which are settings that Pi0 has not encountered during its training.
Figure 6: Total Average success rate across input modalities. For each input modality, we test on 20 trials on each task, then average the results. We split the success rates of seen and unseen instances (detailed in Section IV-E).
IV-C Evaluation Protocol and Metrics
Success Rate: We evaluate each policy on 20 real-world episodes per task and report per-task success split by seen vs. unseen instances (Fig. 6). Our keypoint-based policy consistently outperforms RGB and RGB-D baselines, especially on unseen objects—likely because dense RGB inputs overfit to a single environment. For the UAD baseline, although it can extract dense features from the objects, it fail short with small number of demonstration because of its high-dimension representations. In contrast, the Pi0 foundation model struggled on most tasks. The first reason is that our setup uses only an external camera view, whereas Pi0 was designed to also use a wrist-mounted camera. Secondly, Pi0’s underlying vision-language model (PaLI-GEMMA) is relatively small, which required more data to refine and also made it sensitive to the language instructions.
Figure 7: Semantic Information rubric scores across input modalities on seen instances. For each input modality, we collected 40 trials for each task (we exclude the ball kick task as there is no notion of grasping). The success/failed shown as Y/N in the table for each sub-goal are manually recorded.
Semantic Trajectory Metrics: To diagnose failures beyond binary task success, we create a rubric to score each trajectory (as suggested by  [36]). The rubric breaks captures the success of several intermediate steps and not just the final task success (Fig. 7). In our evaluation, each sub-goal is checked independently, meaning we record whether the policy managed to achieve each individual step. To this end, we found cases where the RGB and RGB-D BC policies attempted to perform the final motion of the task (such as pouring or cutting) before actually grasping the object, indicating a confusion in the sequence of actions. In contrast, our keypoint-based policy typically reaches the final sub-goal, with occasional last-step slippage (object slipping from the gripper), suggesting strong early-stage competence.
Data Scaling: We study the effect of demonstration data size by training each method on progressively smaller subsets of the data. In particular, we downsampled the demonstrations to 50% and 25% of the full set and measured the success rates (Fig. 8. The results show that our keypoint-based method remains highly effective even with far fewer demonstrations, whereas the image-based baselines degrade significantly as the training data is reduced. This highlights the data efficiency of our approach compared to learning directly from raw images.
Figure 8: Afford2Act outperforms all baselines even with just 10 demonstrations, with the performance gap remaining consistent as more demos are added. Standard deviation bars show variability across tasks.
IV-D Keypoints Selections Baselines
After demonstrating that a 2D keypoint-based policy can be highly effective, we ask whether our affordance-driven keypoint selection is essential. We compare our approach with several alternative keypoint selection strategies, each using $K=19$ (15 on object and 4 on end-effector) keypoints and the same downstream policy architecture for fairness:
YOLO [37] + DINO [29]: Use a YOLO object detector to generate bounding boxes, then sample DINO keypoint correspondences within each detected box.
YOLO [37] + Uniform Grid: Uniformly sample a grid of points inside each YOLO-detected bounding box.
DINO [29] Only: Use DINO to generate correspondences over the entire image without bounding-box or filtering.
SAM [38] + DINO: Use the Segment Anything Model (SAM) to obtain segmentation masks, use YOLO to identify the target object’s mask, and then sample DINO correspondences within that segmented region.
Hand-Picked (P3PO): An oracle baseline where an expert manually chooses the keypoints (following P3PO [6]), and these points are then tracked using DIFT + CoTracker.
Fig. 9 summarizes the success rates of these strategies, averaged across all tasks, for both seen and unseen object instances. We find that the methods relying on YOLO detection do not significantly improve generalization to unseen object categories. This is likely because they depend on known object class labels (e.g., always looking for a “cup”), which can be misguiding when the object geometry changes (for example, encountering a teapot that might not be recognized as a cup). It also can suffer in cluttered scenes when the boxes are not tight or when multiple objects are present. In contrast, our affordance-driven approach does not depend on specific object labels; it directly identifies task-relevant regions (interaction hotspots) via the affordance model. We also note that the manually selected keypoints (P3PO) yield a success rate nearly on par with Afford2Act. However, our method achieves this performance without the tedious manual effort of choosing points, and it offers the flexibility of using natural language to specify the task.
IV-E Generalization Analysis
Beyond evaluating overall success, we systematically test four axes of generalization to assess how well Afford2Act transfers to new situations:
  1. 1.
    Unseen Object (Same Category): Test the policy on novel instances of the same category (e.g., a new mug for the “hold” task or a different brush for “brush_with”).
  2. 2.
    Unseen Category: Apply the learned skill to an object from a category not encountered during training (e.g., using a spatula to stir instead of a spoon).
  3. 3.
    Background Variation: Evaluate the policy in environments with different backgrounds or lighting conditions.
  4. 4.
    Distractors: Introduce additional objects or even moving agents (e.g., a person performing an unrelated action nearby) to test the policy’s ability to stay focused on the target task in the presence of distractions.
Fig. 10 illustrates these generalization scenarios and the corresponding performance. Our keypoint-based policy exhibits strong transferability to novel object instances within seen categories and even to objects from entirely new categories (essentially achieving zero-shot generalization). It also remains robust under background changes and in cluttered environments. For example, in one challenging trial, the system successfully transferred the learned affordance to a teapot from a mug in a highly cluttered scene, correctly identifying the meaningful keypoints despite different geometry. Likewise, the policy proved resilient to dynamic distractions, even with a human actively moving nearby.
Figure 9: Total Average success rate across keypoint selection methods.
Figure 10: Top: Average success rate on four scenarios aggregated across all tasks. Bottom: Representative rollouts for each scenario. The small inset labeled “Ref” shows the training reference object for that task.
IV-F Ablation Studies
We conducted additional ablations to highlight the contribution of different components of our pipeline:
Language Prompt Robustness: We varied the wording of the affordance prompt to ensure our approach is not overly sensitive to specific language. Qualitatively, we found that synonyms or related phrases produce similar affordance keypoints. For example, using “sweep” instead of “brush_with” or “pick_up” instead of “hold” resulted in nearly identical affordance predictions on the relevant objects, indicating the model robustly understands the intended action.
Gating Network: We removed the gating network so that each keypoint token receives equal weighting. This modification significantly degraded performance (especially in cluttered scenes), reducing the overall success rate from $86\%$ to $52\%$. During trajectory rollout, the gating mechanism allows the policy to focus only on the keypoints necessary.
Effective Keypoint Count: We analyzed the gating network’s behavior by computing the effective number of keypoints being used by the policy at each time step. Given the gating output weights $w_{i}$} for the $K$ keypoint tokens, we define $K_{\text{eff}}=1/\sum_{i=1}^{K}w_{i}^{2}$. In our experiments, $K_{\text{eff}}$ was typically much smaller than the total number of points $K=19$, as shown in Fig. 11. This implies that the gating network effectively attends to only a subset of the keypoints at any given time, which helps the policy concentrate on the most informative features for the task.
Figure 11: Average effective number of keypoints used (over all time steps) for each task. This data was collected through offline inference on all 40 training demos for each task.
IV-G Long-Horizon Skills via Two-Stage Re-Anchoring
For more complex, multi-phase tasks, we integrate Large Model task planning stage into our framework. We first use a vision-language model (VLM) planner to detect when the first sub-goal of the task has been completed (e.g., when the robot has successfully grasped the object). Once this condition is detected, we re-anchor half of the keypoints by providing an updated affordance prompt for the next phase of the task. In practice, this means replacing roughly half of the original keypoints with new ones tailored to the subsequent sub-task. This re-anchoring allows the policy to adapt to the changed state of the scene and effectively execute the next stage in the trajectory. We experimented with this augmentation on the Stir task, achieving an impressive $90\%$ success rate, further showing promise with incorporating VLMs with affordance-based keypoint selection (see video).
V Conclusions and Future Work
We introduced Afford2Act, a three-stage imitation learning pipeline that leverages semantically meaningful keypoints derived from affordance text prompts. Through six real-world experiments, we demonstrated that Afford2Act is not only data-efficient and performant, but also generalizes effectively to unseen objects, categories, backgrounds, and distractors. While our work represents the first attempt at semantically meaningful keypoint-based imitation learning, the selected keypoints remain fixed during inference and cannot be adapted dynamically to evolving contexts. As discussed in Sec. IV-G, integrating task planning with keypoint resampling is a promising future direction, potentially enabling flexible long-horizon policy execution.

### v2

@@
 Abstract
-Vision-based robot learning often relies on dense image or point-cloud inputs, which are computationally heavy and entangle irrelevant background features.
-Existing keypoint-based approaches can focus on manipulation-centric features and be lightweight, but either depend on manual heuristics or task-coupled selection, limiting scalability and semantic understanding.
-To address this, we propose Afford2Act, an affordance-guided framework that distills a minimal set of semantic 2D keypoints from a text prompt and a single image.
-Afford2Act follows a three-stage pipeline: affordance filtering, category-level keypoint construction, and transformer-based policy learning with embedded gating to reason about the most relevant keypoints, yielding a compact 38-dimensional state policy that can be trained in 15 minutes, which performs well in real-time without proprioception or dense representations.
-Across diverse real-world manipulation tasks, Afford2Act consistently improves data efficiency, achieving an 82% success rate on unseen objects, novel categories, backgrounds, and distractors.
-Visualizations available at: https://afford2act.github.io/
+Training vision-based manipulation policies that generalize beyond the training scenes remains challenging because dense visual observations entangle task-relevant cues with nuisance variation such as background, lighting, and instance-specific appearance.
+Keypoint representations offer a compact alternative, but automatically selecting task-relevant keypoints that transfer across objects and instructions is non-trivial.
+We present Afford2Act (Affordance-Guided Keypoints to Act), which uses language-conditioned affordance priors to localize where interaction should occur and extract a small set of semantic 2D keypoints that capture functionally relevant object parts.
+Since keypoint relevance can change over the course of an interaction, we use a lightweight attention-and-gating policy to prioritize keypoints for control.
+Across six real-robot manipulation tasks, Afford2Act achieves 82% success on held-out objects and is robust to clutter, distractors, and open-vocabulary prompt variations.
 I Introduction
-Learning effective robot manipulation policies hinges on overcoming a fundamental representation challenge: how to extract the manipulation-related features which are both compact and expressive?
-Dense visual inputs such as images [1], pointclouds [2], or semantic dense features [3, 4] provide rich details, but burden a policy with extraneous information.
-In contrast, 2D keypoint [5] offers a sparse state representation that highlights only the essential features of objects’ motions while still being representative enough to learn a good policy.
-This sparsity directly enables a lightweight policy with far fewer but important features than dense representation and support effective and real-time inference with less computational demand.
-However, the real challenge is in discovering these keypoints automatically and reliably: picking too few or irrelevant points can omit critical information, while relying on human insight for what points to use does not scale.
-In this paper, we present a method for identifying task-relevant object points that are both minimal (compact) and sufficient (expressive) for effective robot manipulation.
-The problem of identifying keypoints has received a lot of attention recently.
-Some previous work utilizes manual selection [6] and prompts large Vision-Language Models (VLMs) to mark affordances on images [7, 3].
-While these methods can provide useful cues, they depend on human guidance or external knowledge, making them labor-intensive and domain-specific.
-At the other end of the spectrum, some approaches forgo sparsity altogether in favor of dense feature grid-based representations.
-For instance, all previous work [8, 9, 10, 11] learn from dense grid of points in the scene or the object.
-While dense features are expressive, they are not compact and incur heavy computational demands.
-Furthermore, dense features may also entangle irrelevant dynamics (such as background features), which limits scalability and focus.
-Figure 1: Top: policy trained with original prompts (“hold”, “stir”) executes the actions over time ($t$).
-Middle: at test time, the same policy generalizes to synonym prompts (“pick”, “mix”).
-Bottom: affordance-guided keypoints with learned importance (green; variable count; ring radius $\propto$ attention) remain stable under unseen object shapes, lighting conditions, dynamic distractors, and scene clutter—leading to successful rollouts.
+Learning effective robot manipulation policies hinges on overcoming a fundamental representation challenge: how can we extract manipulation-related features that are both compact and expressive?
+Dense visual inputs such as images [4], point clouds [35], or semantic features [26, 30] provide rich detail, but may burden a policy with extraneous information [27].
+In contrast, 2D keypoints [7, 20] offer a sparse input representation that highlights only the essential features of objects while still being representative enough to learn a good policy.
+This sparsity directly enables a lightweight policy with far fewer, but more important, features than dense representations and supports effective real-time inference with lower computational demand.
+However, the real challenge is discovering these keypoints automatically and reliably: picking too few or irrelevant points can omit critical information, while relying on human insight for point selection does not scale.
+Automatic keypoint discovery for manipulation has recently gained momentum, with most methods falling into three paradigms.
+Motion-driven approaches identify keypoints as pixels (or point tracks) that move significantly when an object is manipulated, and then learn control from the resulting point tracks [2, 32, 34].
+While effective when object motion is clean and informative, they can degrade in non-static scenes where background motion introduces spurious tracks, making it harder for the policy to remain reliable under dynamic distractors (e.g., a person operating in the same space as the manipulator).
+A second family of methods learns keypoints jointly with the policy, optimizing point selection end-to-end for task performance [36].
+This can produce compact representations, but often couples the discovered points to the training data, making transfer to new objects or task intents less straightforward.
+Finally, a complementary direction seeks object-centric keypoints by extracting by grouping the semantic keypoints [29, 8].
+These points can be helpful for consistent localization across frames and instances, but they are often agnostic to interaction: they describe where the object is and how it appears, rather than where the robot should act.
+Consequently, such keypoints may remain largely unchanged even when the intended use of the same object changes.
+In manipulation, however, relevance is inherently task-conditioned: the same tool supports multiple feasible interactions, and different instructions should highlight different regions (e.g., a brush can be grasped to lift/hold, or oriented and contacted to sweep; Fig.
+4).
+This motivates an open question: can we automatically extract functionally grounded, task-conditioned keypoints that generalize across object shapes while remaining robust to distractors?
+Figure 1: Top: Visualization of affordance-guided keypoint extraction with an example language prompt.
+Our pipeline reduces a $224\times 224$ image to 19 keypoints (38 dimensions).
+There are up to 15 keypoints on the object’s actionable region and 4 on the robot end-effector.
+The small sub-box in the top-right image is the affordance heatmap.
+Bottom: Examples of unseen instances to which our policy generalizes.
 Figure 2: Overview of the Afford2Act pipeline.
-We extract task-relevant object keypoints using affordance masks and DINO filtering, track them across frames, and embed them with a transformer and attention gating.
-The policy head predicts pose and gripper commands from these compact keypoint representations, enabling vision-based manipulation without requiring proprioception.
-Recently, there has been work on automatically selecting a subset of keypoints from a dense grid [12, 13, 14] in an end-to-end manner.
-In particular, ATK [12] jointly optimizes a masking model for selecting a compact keypoint set and a task-specific control policy, yielding strong performance on challenging visuomotor tasks.
-However, there are two key limitations in ATK that prohibit it from generalizing: (i) Joint training couples the masking model to the task, so extending to a new task requires retraining both components (e.g., moving from mug picking from left to right); (ii) Existing evaluations emphasize robustness to positional variation and visual perturbations, whereas other generalization regimes, duplicate objects and cross-category transfer (e.g., from mugs to teapots), have received limited attention.
-In summary, prior works either require significant prior knowledge (manual keypoints), incur high computational overhead (dense grids), or struggle with stability and semantic generalization.
-This state of affairs motivates a core question for robot learning: How can a robot learn to identify the most relevant interaction points, the task hotspots, directly from visual inputs?
-To address this challenge, we propose Afford2Act: Affordance-Guided Keypoints to Act, which uses text-prompted affordance cues to automatically select task-relevant keypoints, with no manual labels.
-This pipeline contains three stages: (i) Affordance Filtering (e.g., “hold” $\rightarrow$ mug handle), (ii) Keypoint Pool Construction from a small set of masked DINO anchors, and (iii) Keypoint-Based Policy Learning, where a transformer-based encoder followed by a gating mechanism learns the importance of each keypoint.
-Because mask support and geometry vary, the number of selected keypoints per object also varies; thus, correspondence is defined as a flexible set of affordance-region points.
-As visualized in Fig.
-1, the policy typically attends to a small effective subset (Fig.
-11) of keypoints.
-With a 38-dimensional input, policies remain robust under open-vocabulary prompts and scene changes across six real-world tasks.
-Evaluating across six real-world tasks, Afford2Act consistently outperforms keypoint-selection baselines and RGB/RGB-D policies, and stays robust and generalizable to synonym prompts, lighting, distractors, and clutter (Fig.
-1).
+Step 1 takes single-view RGB demonstrations and language prompts as input.
+Step 2 performs affordance-guided keypoint distillation by localizing the actionable region and selecting reference anchor keypoints with DINO features.
+Step 3 establishes cross-instance correspondence to preserve semantic keypoint consistency across object instances.
+Step 4 tracks keypoints over time and learns a compact transformer-based policy with gating to prioritize task-relevant tokens when predicting end-effector motion and gripper commands.
+To address this challenge, we turn to affordances [26, 18].
+A natural way to make feasible interactions explicit is to link an instruction (e.g., hold, cut, pour) to actionable regions in the scene.
+Building on this intuition, we introduce AFFORD2ACT: Affordance-Guided Keypoints to Act.
+Our approach first filters candidate points to instruction-relevant regions, then preserves functional keypoint identity across object instances through correspondence, and finally adapts keypoint importance during control with gating that dynamically re-weights keypoints over time, focusing on different interaction points at different stages of execution.
+Together, these components produce a compact state tied to functionally meaningful object parts rather than appearance, enabling efficient policy learning from a single external RGB view.
+Across six real-world manipulation tasks, this representation remains robust to novel objects, clutter, and open-vocabulary prompts.
 Contributions. We present:
-  1. 1.
-    A unified Affordance-Guided Keypoint Distillation framework that extracts sparse, semantically meaningful 2D keypoints directly from text prompts.
-  2. 2.
-    An efficient, attention-based framework that trains a compact keypoint-guided policy in $\sim$15 minutes.
-  3. 3.
-    Extensive real-robot experiments in six tasks, demonstrating superior 82% success in unseen instances and robustness to open-vocabulary prompts (Fig. 1).
-Figure 3: Real-world rollouts of Afford2Act on Pour (top) and Cut (bottom) tasks.
-In each sequence (left to right: approach $\rightarrow$ interaction $\rightarrow$ completion), green dots mark keypoints automatically selected on the affordance region (e.g., knife for “cut”, handle/spout for “pour”).
-The translucent ring size reflects the attention weight at each step.
-While keypoints remain fixed on task-relevant parts, the policy dynamically shifts its focus: for “cut”, the attention moves from the blade tip toward the handle; for “pour”, from the edge to the main handle.
-This adaptive attention enables robust, context-aware manipulation.
+  • A language-conditioned affordance-guided keypoint extraction pipeline that selects sparse 2D keypoints from interaction-relevant regions without manual keypoint specification (Sec. III-A).
+  • A correspondence-and-gating policy formulation that preserves keypoint consistency across object instances while adapting keypoint relevance over the course of manipulation (Sec. III-B).
+  • Extensive real-robot evaluation on six tasks, showing robust transfer to held-out objects, scene variations, and prompt paraphrases with a compact keypoint state (Secs. IV, IV-E; Fig. 1).
+Figure 3: Real-world rollouts of Afford2Act on unseen instances.
+Rows show Pour (top), Cut (middle), and Stir (bottom), with each sequence progressing left-to-right from approach $\rightarrow$ interaction $\rightarrow$ completion.
+Green dots denote affordance-selected keypoints, and translucent ring size indicates attention weight over time.
+Keypoints remain anchored to task-relevant parts, while attention shifts with phase (e.g., cut: blade tip $\rightarrow$ handle; pour: edge/spout $\rightarrow$ main handle; stir: handle $\rightarrow$ bowl), enabling robust, context-aware manipulation.
 II Related Work
 Visual Representations for Policy Learning: Choosing the right visual representation is central to building robust manipulation policies.
-Classic approaches rely on raw sensory data which includes images, depth maps, or point clouds to provide a complete view of the environment [1, 2, 15].
-While these representations are rich, they also saddle policies with high-dimensional inputs, requiring heavy data regimes and often resulting in policies that overfit to appearance rather than task.
+Classic approaches rely on raw sensory inputs such as images, depth maps, or point clouds to provide a complete view of the environment [35].
+While these representations are rich, they also saddle policies with high-dimensional inputs, requiring large data regimes and often resulting in policies that overfit to appearance rather than task.
 As a remedy, manipulation-centric and object-centric representations [26, 13, 9, 24, 33, 25, 12] filter out much of the visual clutter, focusing policy learning on the handful of features that matter most for action.
-However, these strategies often rely on hand-tuned detectors or dense input from images 3D reconstructions, limiting scalability and flexibility.
-Our method bypasses these constraints, extracting a compact, affordance-driven set of 2D keypoints that serve as a focused visual state, requiring neither dense geometry nor extensive supervision.
+However, these strategies often rely on hand-tuned detectors or dense inputs from images or 3D reconstructions, limiting scalability and flexibility.
+Our method bypasses these constraints by extracting a compact, affordance-driven set of 2D keypoints that serves as a focused visual state, requiring neither dense geometry nor extensive supervision.
 Keypoints and Automatic Anchor Discovery: Keypoints, when properly chosen, offer a natural low-dimensional representation that encodes an object’s essential affordances [17, 32, 34, 29].
-Earlier work often assumed keypoints could be specified by experts, or required laborious labeling to maintain semantic consistency across objects [6, 22].
-More recent efforts automate this selection, using object detector [10, 11] or learning which points are most predictive for the task at hand [12, 8].
-For keypoint sampling from detected objects, the representations are still dense and lack semantic meaning.
+Earlier work often assumed keypoints could be specified by experts or required laborious labeling to maintain semantic consistency across objects [17, 20].
+More recent efforts automate this selection, using object detectors [32, 34] or learning which points are most predictive for the task at hand [36, 31].
+When keypoints are sampled from detected objects, the resulting representations are still dense and often lack semantic meaning.
 ATK, for instance, integrates keypoint selection with policy optimization, but can produce unstable and semantically ambiguous anchors.
@@
 Overview.
-Given single-view demonstrations $\mathcal{D}=\{(I^{(d)}_{1:T_{d}},\,\mathbf{u}^{(d)}_{1:T_{d}},\,A^{(d)})\}_{d=1}^{D}$, where $I^{(d)}_{t}\in\mathbb{R}^{H\times W\times 3}$ are images, $\mathbf{u}^{(d)}_{t}\in\mathbb{R}^{m}$ are actions (e.g., 6-DoF deltas and a gripper scalar), and $A^{(d)}$ is a text prompt, Afford2Act: (i) localizes the actionable region via a prompt-conditioned affordance mask (Section III-A), (ii) selects a small set of semantic keypoints on the reference frame (Section III-A), (iii) matches them within the mask on new instances to establish flexible correspondence ($\leq 15$) for various geometry and lightening (Section III-A), (iv) jointly tracks selected keypoints and encodes them into a transformer with a gating layer that learns phase-adaptive weights to encode a compact scene embedding for policy learning (Section III-B).
+Given single-view demonstrations $\mathcal{D}=\{(I^{(d)}_{1:T_{d}},\,\mathbf{u}^{(d)}_{1:T_{d}},\,A^{(d)})\}_{d=1}^{D}$, where $I^{(d)}_{t}\in\mathbb{R}^{H\times W\times 3}$ are images, $\mathbf{u}^{(d)}_{t}\in\mathbb{R}^{m}$ are actions (e.g., 6-DoF deltas and a gripper scalar), $T_{d}$ is the timestep at the end of the demonstration, and $A^{(d)}$ is an affordance text prompt, Afford2Act: (i) localizes the actionable region via a prompt-conditioned affordance mask (Section III-A), (ii) selects a small set of semantic keypoints on the reference frame (Section III-A), (iii) matches them within the mask on new instances to establish flexible correspondence with up to 15 keypoints across variations in geometry and lighting (Section III-A), and (iv) jointly tracks selected keypoints and encodes them into a transformer with a gating layer that learns phase-adaptive weights to produce a compact scene embedding for policy learning (Section III-B).
 The overall framework is shown in Fig.
 2.
-Formal Setup.
-Each demonstration $d$ is summarized by trajectories of $K$ 2D keypoints $\mathbf{P}^{(d)}=\{(x^{(d)}_{t,i},\,y^{(d)}_{t,i})\}_{t=1:T_{d},\;i=1:K}$, where indices $i$ refer to functional parts within the affordance region (e.g., “handle tip,” “blade edge”).
-This sparse, semantic representation lets the policy reason over roles rather than raw pixels.
+Setup.
+Each demonstration $d$ is summarized by trajectories of $K$ 2D keypoints, represented as $\mathbf{P}^{(d)}=\{(x^{(d)}_{t,i},\,y^{(d)}_{t,i})\}_{t=1:T_{d},\;i=1:K}$, where indices $i$ refer to functional parts within the affordance region (e.g., “handle tip,” “blade edge”).
+This sparse semantic representation lets the policy reason over roles rather than raw pixels.
 III-A Affordance-Guided Keypoint Representation
 Our pipeline for extracting these keypoints proceeds in four stages, all fully automatic and requiring no manual annotation (see Fig. 2).
-Affordance Localization: Given the first frame $I_{1}$ of a demonstration affordance prompt, we use One-Shot Open Affordance model  [28] $f_{\text{aff}}$ to produce a heatmap: $H=f_{\text{aff}}(I_{1},A)\in[0,1]^{H\times W}$.
-This heatmap indicates regions of the scene that likely correspond to the intended action.
-To focus only on the actionable surface, we apply a robust quantile threshold (along with mild smoothing) to obtain a binary mask: $M(x,y)=\mathbb{1}\big[H(x,y)\geq\tau_{q}(H)\big]$.
+Affordance Localization: Given the first frame $I_{1}$ of a demonstration and an affordance prompt $A$, we use a pre-trained One-Shot Open Affordance model [18] $f_{\text{aff}}$ to produce a heatmap: $H=f_{\text{aff}}(I_{1},A)\in[0,1]^{H\times W}$.
+This heatmap indicates regions of the scene that correspond to the intended action.
+To focus only on the actionable surface, we apply a robust quantile threshold (along with mild smoothing) to obtain a binary mask: $M(x,y)=\mathbb{1}\big[H(x,y)\geq\tau_{q}(H)\big]$, where $\tau_{q}(H)$ is the threshold value.
 This mask captures fine structures such as a knife blade or cup rim and robustly filters out clutter.
-Text-based prompts provide a strong semantic prior, allowing the same model to adapt affordance detection across novel tools or object categories with minimal (such as the examples in Fig.
-3).
-Reference Keypoint Extraction: On a reference frame (in our case, the first frame of the first demonstration), $I_{1}^{r}$, we compute dense feature maps using a Vision Transformer backbone (ViT/DINO) [29]: $F_{r}=\phi(I^{r}_{1})\in\mathbb{R}^{H\times W\times D}$.
-We then restrict attention to the ROI defined by the affordance mask $M$ and cluster the features within that region to obtain $N$ reference anchor points:
-$\mathcal{A}_{r}=\left\{(p_{i}^{r},\,\mathbf{f}_{i}^{r})\right\}_{i=1}^{N},\qquad p_{i}^{r}\in\Omega(M),\ \mathbf{f}_{i}^{r}=F_{r}(p_{i}^{r})$
-This step is fully automatic and scales to new categories, in contrast to manual annotation-based approaches [6, 7].
-Cross-instance Correspondence: Given a new demonstration, we extract features from its first frame, $F_{t}=\phi(I^{t}_{1})$.
-From each keypoint, $(p_{i}^{r},\mathbf{f}_{i}^{r})$ in the reference anchor, $\mathcal{A}_{r}$, we compute a cosine similarity map restricted to the affordance ROI in the new frame:
-$S_{i}(x,y)=\left\langle\frac{\mathbf{f}_{i}^{r}}{\|\mathbf{f}_{i}^{r}\|},\,\frac{F_{t}(x,y)}{\|F_{t}(x,y)\|}\right\rangle,\quad(x,y)\in\Omega(M_{t})$
-We assign the keypoint with the highest similarity,
-$p^{t}_{i}=\arg\max_{(x,y)\in\Omega(M_{t})}S_{i}(x,y),\quad c^{t}_{i}=\max_{(x,y)\in\Omega(M_{t})}S_{i}(x,y)$
-This matching ensures that the index $i$ refers to the same functional part (e.g., “handle tip”) across all demonstrations, even as object geometry and appearance vary.
-By constraining matches to the affordance mask, we avoid drifting to irrelevant but visually similar regions.
-Temporal Keypoint Tracking: Once all $K$ keypoints are initialized on the first frame, we track them across the demonstration using a transformer-based tracker such as CoTracker [30]: $\{p_{t,i}\}_{t=1:T,\,i=1:K}=g_{\text{trk}}(I_{1:T},\,\{p^{1}_{i}\}_{i=1}^{K})$
-Joint tracking improves robustness to occlusion and viewpoint changes, leveraging correlations among keypoints.
-The output is a set of trajectories in a semantic space where each index $i$ always refers to the same actionable region.
-By the end of this process, every demonstration is described by a trajectory of $K$ task-relevant, affordance-grounded keypoints.
-This sparse and semantically aligned representation serves as the input to our downstream policy (Section III-B), bridging the gap between the efficiency of keypoints and the semantic expressiveness of affordances, without requiring heavy 3D pipelines or manual keypoint curation.
+Reference Keypoint Extraction: On a reference frame (the first frame of the first demonstration), $I_{1}^{r}$, we compute dense feature maps using a DINO backbone ($\phi$[22]: $F_{r}=\phi(I^{r}_{1})\in\mathbb{R}^{H\times W\times D}$.
+We use DINO features because they provide a robust, part-sensitive similarity metric that enables more reliable cross-instance correspondence.
+We then restrict attention to the ROI  [^1] ROI stands for Region of Interest, i.e., the cropped region of the affordance mask.
+defined by the affordance mask $M$ and cluster the features within that region to obtain $K$ reference anchor points in a reference anchor set $\mathcal{A}_{r}$:
+$\mathcal{A}_{r}=\left\{(p_{i}^{r},\,\mathbf{f}_{i}^{r})\right\}_{i=1}^{K},\qquad p_{i}^{r}\in\Omega(M),\ \mathbf{f}_{i}^{r}=F_{r}(p_{i}^{r})$
+where $\Omega(M)$ is the set of pixel coordinates covered by the mask $M$. This step is fully automatic and scales to new categories, in contrast to manual annotation-based approaches [17, 19].
+Cross-instance Correspondence: Given a new demonstration, we compute a dense feature map from its first frame, $F_{t}=\phi(I^{t}_{1})$.
+For each reference anchor keypoint $(p_{i}^{r},\mathbf{f}_{i}^{r})\in\mathcal{A}_{r}$, we match it in the new frame by evaluating cosine similarity between the reference descriptor and the per-pixel features, restricted to the affordance ROI:
+$S_{i}(x,y)=\left\langle\frac{\mathbf{f}_{i}^{r}}{\|\mathbf{f}_{i}^{r}\|},\,\frac{F_{t}(x,y)}{\|F_{t}(x,y)\|}\right\rangle,\quad(x,y)\in\Omega(M_{t}).$
+We then select the best-matching location and its confidence as:
+$p^{t}_{i}=\arg\max_{(x,y)\in\Omega(M_{t})}S_{i}(x,y),\quad c^{t}_{i}=\max_{(x,y)\in\Omega(M_{t})}S_{i}(x,y).$
+This matching ensures that the index $i$ refers to the same functional part (e.g., “handle tip”) across demonstrations, even as object geometry and appearance vary.
+By constraining matches to the affordance mask, we reduce spurious matches to visually similar but task-irrelevant regions.
+Temporal Keypoint Tracking: After initializing all $K$ keypoints in the first frame, we track them over the demonstration using a transformer-based point tracker (e.g., CoTracker [14]).
+This produces a temporally ordered set of keypoint locations,
+$\{(p_{1,1},\ldots,p_{1,K}),\;(p_{2,1},\ldots,p_{2,K}),\;\ldots,\;(p_{T_{d},1},\ldots,p_{T_{d},K})\},$
+where $p_{t,i}$ denotes the 2D location of keypoint $i$ at time $t$ and $T_{d}$ is the demonstration length.
+Joint keypoint tracking with Co-Tracker improves robustness to occlusion (e.g., when the gripper blocks the view of the handle during grasping) by exploiting correlations among keypoints, and preserves index consistency so that each $i$ continues to refer to the same actionable region over time.
+After this step, each demonstration is represented as a trajectory of $K$ task-relevant, instruction-conditioned keypoints.
+We use this sparse, temporally consistent representation as input to the downstream policy (Section III-B), without requiring dense 3D reconstruction or manual keypoint specification.
 III-B Policy Design
-At each interfere step $t$, taking the $K$ keypoints from keypoint distillation $\{p_{t,i}\}_{i=1}^{K}$, the question becomes: how do we architect a policy, $\pi_{\theta}$, that prioritizes the most task-relevant keypoints?
-After embedding each keypoint, $z^{1}=[\phi_{\theta}(p_{1})\dots\phi_{\theta}(p_{K})]^{\top}$, we then compute contextualized token features using a Transformer chunk, $z=\mathcal{T}_{\theta}(z_{1})$, similar to [31, 32] so that each keypoint is interpreted in the context of the others (e.g., a keypoint on the blade of a knife carries a different affordance than one on the handle).
-Since the number of keypoints $m$ can exceed the task-relevant cardinality, we introduce a gating network, $g_{\psi}$, that predicts a non-negative importance weight for each keypoint before pooling them into a single scene feature vector $h$:
-$\hat{w}=g_{\psi}(z)=\operatorname{softmax}(w_{g}^{\top}z+b_{g}),\hat{z}=\hat{w}^{\top}z,\quad h=\frac{1}{K}\sum_{i=1}^{K}\hat{z}_{i}$
-Then, the scene embedding, $h$ is passed through a shared action head that is later split into two linear layers: one for the delta actions and the other for the gripper state.
-While we use a two-layer MLP, the head is modular and can be replaced by a generative policy head [33, 34, 1].
+At each inference step $t$, we are given the $K$ distilled keypoints $\{p_{t,i}\}_{i=1}^{K}$.
+The policy must transform this unordered set into an action while prioritizing the keypoints that are most informative for the instructed manipulation.
+We first embed each keypoint into a token using a learnable embedding function $\psi_{\theta}$, forming a token set
+$Z_{t}=[z_{t,1},\ldots,z_{t,K}]^{\top},\qquad z_{t,i}=\psi_{\theta}(p_{t,i}).$
+We then compute contextualized token features with a Transformer encoder block: $\tilde{Z}_{t}=\mathcal{T}_{\theta}(Z_{t})$, so that each keypoint is interpreted in the context of the others (e.g., a point on a knife blade should be treated differently from a point on the handle) [11, 37].
+Gating network.
+Keypoint distillation may return redundant or noisy candidates within the actionable region.
+We therefore introduce a gating network $g_{\theta}$ that assigns a non-negative importance weight to each token, allowing the policy to down-weight task-irrelevant keypoints.
+This explicitly separates candidate generation from relevance estimation within the control policy.
+Concretely, we compute per-keypoint scores and normalize them with a softmax:
+$s_{t,i}=w_{g}^{\top}\tilde{z}_{t,i}+b_{g},\qquad\hat{w}_{t}=\operatorname{softmax}([s_{t,1},\ldots,s_{t,K}]),$
+where $\hat{w}_{t}\in\mathbb{R}^{K}$ and $\hat{w}_{t,i}\geq 0$ with $\sum_{i=1}^{K}\hat{w}_{t,i}=1$.
+We form gated token features and aggregate them into a compact scene embedding: $h_{t}=\sum_{i=1}^{K}\hat{w}_{t,i}\,\tilde{z}_{t,i}$.
+The resulting $h_{t}$ is then passed through a shared MLP action head, which branches into two linear outputs: one predicts the end-effector delta action and the other predicts the gripper command.
 IV Experiments
+Figure 4: (Left) Our physical setup, consisting of a UR3e robotic arm and a side-mounted RealSense D435i camera. (Right) Objects used in seen (right) and unseen (left) scenarios.
+Figure 5: Total average success rate across keypoint selection methods.
+Figure 6: Qualitative generalization on two tasks.
+Rows show rollouts for Pour (top) and Kick (bottom).
+Left: seen training objects.
+Right: held-out settings with unseen scenes (new backgrounds/tabletops) and, for Pour, unseen object instances.
+Each panel shows three key frames (setup, approach, completion).
+Without test-time adaptation, Afford2Act transfers by focusing on functional parts (mug handle/ball) while ignoring clutter.
 Our experiments are designed to evaluate three claims:
   1. 1.
-    Keypoint Effectiveness: Can a policy trained exclusively on a compact set of 2D keypoints outperform policies trained on full images, dense features, and even a large vision-language foundation model?
+    Keypoint Priors (E1): How does our method compare against alternative keypoint selection strategies, and can our fully automatic pipeline match the performance of manually selected keypoints?
+(See Sec.
+IV-C.)
   2. 2.
-    Keypoint Priors: Which alternative keypoint selection strategies (e.g., using object bounding box, uniform grids of points, or manually selected points) can match the informativeness of our affordance-guided keypoints?
+    Keypoint Effectiveness (E2): Can a policy trained exclusively on a compact set of 2D keypoints outperform policies trained on full images or dense affordance features? (See Sec. IV-D.)
   3. 3.
-    Generalization: How well does our policy generalize to new scenarios, including novel objects within seen categories, entirely new object categories, the presence of distractor objects, and other scene variations?
-Additionally, we perform targeted ablation studies to test the robustness to different language prompts, to verify the necessity of the gating network, and to explore an extended two-stage pipeline for long-horizon tasks.
-Figure 4: (Left) A figure of our physical setup, composing of a UR3e Robotic arm and one side RealSenseD435i Camera. (Right) shows the objects we used for seen (right) and unseen (left) scenarios.
-Figure 5: Qualitative generalization on two tasks.
-Each row shows rollouts for Pour (top) and Kick (bottom).
-Left: training objects (seen).
-Right: held-out settings with unseen scenes (new backgrounds/tabletops) and, for Pour, unseen object instances.
-Within each panel we show three key frames (setup, approach, completion).
-Without any test-time adaptation, Afford2Act trained from single-view RGB using affordance-guided keypoints successfully transfers by attending to the functional part (mug handle / ball) while ignoring clutter.
+    Generalization (E3): How well does our policy generalize to new scenarios, including novel objects within seen categories, new object categories (sharing the same functionality), the presence of distractor objects, and other scene variations?
+(See Sec.
+IV-E.)
+Additionally, we perform targeted ablation studies to test robustness to different language prompts, verify the necessity of the gating network, and explore an extended two-stage pipeline for long-horizon tasks.
 IV-A Setup and Tasks
-Our method is driven by task semantics expressed as natural-language prompts that condition an affordance model.
-We focus on these core affordances—hold, cut_with, brush_with, pour, and kick—and systematically construct a variety of manipulation tasks by pairing each affordance with diverse objects and contexts.
-The mapping from affordances to example tasks is as follows:
-Hold: Grasp and lift an object with a handle (e.g., a mug) using the hold affordance. (Fig. 1)
-Cut: Perform a cutting motion on an object using a knife, guided by the cut_with affordance. (Fig. 3)
-Brush: Sweep debris using a hand brush. This task leverages both the hold and brush_with affordances in coordination (holding the brush and using it to brush). (Fig. 2)
-Pour: Execute a pouring motion with a vessel (e.g., tipping a cup or pitcher) using the pour affordance. (Fig. 5)
-Stir: Stir the contents of a bowl using a spoon or similar tool, guided by a combination of hold (to grasp the spoon) and a stir affordance. (Fig. 1)
-Kick: Push a ball into a goal using the kick affordance. 5
-For each task, we train a policy and test it with a variety of seen and novel objects (see Fig.
+Our evaluation is organized around natural-language prompts that specify the intended interaction.
+We focus on a small set of affordances (hold, cut_with, sweep, pour, stir, kick) and instantiate them into six manipulation tasks: grasp-and-lift, cutting, brushing/sweeping (combining hold and sweep), pouring, stirring (combining hold and stir), and ball-kicking (Fig.
 4).
-The objects are physically diverse in size, shape, appearance, and material.
-This diversity allows us to evaluate generalization at both the instance level (new objects of a trained category) and the category level (objects from entirely new categories).
-We collected 40 teleoperated demonstrations per task for training, with object positions randomized across demonstrations.
-IV-B Input Modality Baselines
-We benchmark our approach against four baseline methods to evaluate the keypoints effectiveness:
+For each task, we train an independent policy and evaluate on both seen and novel objects.
+The object sets vary widely in size, geometry, appearance, and material, allowing us to measure transfer to new instances as well as to novel categories that share similar functional parts (Fig.
+4).
+IV-B Training and Evaluation Protocol
+We train one policy per task using 40 teleoperated demonstrations and evaluate success on 20 real-world episodes per task, reported separately for seen and unseen objects.
+For keypoint-selection comparisons, all methods share the same downstream policy architecture and a fixed keypoint count of $K=19$ (15 object keypoints and 4 end-effector keypoints), so differences reflect representation quality rather than controller capacity.
+For semantic trajectory analysis, we annotate 40 seen-object trials per task, excluding ball-kicking because it has no grasp phase, and for data-scaling experiments, each method is retrained on reduced demonstration subsets.
+This efficiency comes from running pretrained affordance and correspondence modules only on the first frame, then training the policy on sparse 2D keypoint trajectories.
+In our implementation, training one Afford2Act policy from 40 demonstrations takes approximately 15 minutes on a single RTX 2080 GPU.
+IV-C Keypoint Priors (E1): Baseline Comparisons
+To demonstrate the strength of affordance-guided keypoint selection, we compare our approach with several alternative keypoint selection strategies, each using $K=19$ (15 on the object and 4 on the end-effector) keypoints and the same downstream policy architecture for fairness:
+YOLO [23] + DINO [22]: Use a YOLO object detector to generate bounding boxes, then sample DINO keypoint correspondences within each detected box.
+YOLO [23] + Uniform Grid: Uniformly sample a grid of points inside each YOLO-detected bounding box.
+DINO [22] Only: Use DINO to generate correspondences over the entire image without bounding-box or filtering.
+SAM [15] + DINO: Use the Segment Anything Model (SAM) to obtain segmentation masks, then use YOLO to identify the target object’s mask and sample DINO correspondences within that segmented region [29].
+Hand-Picked (P3PO-style): A manual expert-selected baseline where keypoints are specified per task (following P3PO [17]) and then tracked using DIFT + CoTracker.
+Fig.
+5 summarizes the success rates of these strategies, averaged across all tasks, for both seen and unseen object instances.
+We find that methods relying on YOLO detection do not significantly improve generalization to unseen object categories.
+This is likely because they depend on known object class labels (e.g., always looking for a “cup”), which can be misleading when object geometry changes (e.g., encountering a teapot that might not be recognized as a cup).
+They can also suffer in cluttered scenes when boxes are not tight or when multiple objects are present.
+In contrast, our affordance-driven approach does not depend on specific object labels; it directly identifies task-relevant regions (interaction hotspots) via the affordance model.
+The manual expert-selected baseline serves as an upper-bound reference for semantically meaningful sparse keypoints.
+Afford2Act reaches comparable performance without per-task point specification, replacing manual geometric priors with a fully automatic, language-conditioned procedure.
+IV-D Keypoint Effectiveness (E2): Input Modality Baselines
+To showcase the strength of keypoints over dense inputs, we conduct an extensive benchmark of our approach against three baselines across success rate, semantic trajectory metrics, and data scaling:
+  • UAD-2D-BC: The same framework as RGB-BC, but the input images are replaced with affordance masks generated by UAD [26].
   • RGB-BC: A behavior cloning policy trained on raw RGB images, using a ResNet-18 visual encoder and a multilayer perceptron (MLP) for the action output.
-  • RGB-D-BC: The framework as RGB-BC, but trained on 4-channel RGB-D images (RGB plus monocular depth from an Intel RealSense camera).
-  • UAD-2D-BC: The same framework as RGB-BC, but the images are replaced with the affordance mask generated by UAD [3].
-This baseline can verify the effectiveness of sparse keypoints rather than dense representations.
-  • Foundation Model (Pi0): A large vision-language-action foundation model (Pi0) [35] fine-tuned on our demonstration data.
-We evaluate this model with a single side camera on a UR3e robot arm, which are settings that Pi0 has not encountered during its training.
+  • RGB-D-BC: The same framework as RGB-BC, but trained on 4-channel RGB-D images (RGB plus monocular depth from an Intel RealSense camera).
+We also evaluated the Pi0 foundation model [3], but did not obtain any successful rollouts in our setting.
+This result is expected given the mismatch in assumptions: our experiments use a single external camera, whereas Pi0 is designed to leverage additional wrist-mounted viewpoints.
+In addition, our robot’s embodiment (UR3e) is not represented in Pi0’s training distribution [28].
 Figure 7: Total average success rate across input modalities.
-For each input modality, we test on 20 trials on each task, then average the results.
-We split the success rates of seen and unseen instances (detailed in Section IV-E).
-IV-C Evaluation Protocol and Metrics
-Success Rate: We evaluate each policy on 20 real-world episodes per task and report per-task success split by seen vs.
+For each input modality, we run 20 trials per task and then average the results.
+We report seen and unseen performance separately (detailed in Section IV-E).
+Evaluation Metrics.
+Success Rate: We evaluate each policy on 20 real-world episodes per task and report per-task success rates split by seen vs.
 unseen instances (Fig.
-6).
-Our keypoint-based policy consistently outperforms RGB and RGB-D baselines, especially on unseen objects—likely because dense RGB inputs overfit to a single environment.
-For the UAD baseline, although it can extract dense features from the objects, it fail short with small number of demonstration because of its high-dimension representations.
-In contrast, the Pi0 foundation model struggled on most tasks.
-The first reason is that our setup uses only an external camera view, whereas Pi0 was designed to also use a wrist-mounted camera.
-Secondly, Pi0’s underlying vision-language model (PaLI-GEMMA) is relatively small, which required more data to refine and also made it sensitive to the language instructions.
+7).
+Our keypoint-based policy consistently outperforms RGB and RGB-D baselines, especially on unseen objects, because dense pixel inputs tend to overfit to a single environment.
+For the UAD baseline, although it can extract dense features from objects, it falls short with a small number of demonstrations because of its high-dimensional representations.
 Figure 8: Semantic information rubric scores across input modalities on seen instances.
-For each input modality, we collected 40 trials for each task (we exclude the ball kick task as there is no notion of grasping).
-The success/failed shown as Y/N in the table for each sub-goal are manually recorded.
+For each input modality, we collected 40 trials per task (excluding ball-kicking, which has no grasping sub-goal).
+Success/failure for each sub-goal is shown as Y/N and was recorded manually.
 Semantic Trajectory Metrics: To diagnose failures beyond binary task success, we create a rubric to score each trajectory (as suggested by [16]).
-The rubric breaks captures the success of several intermediate steps and not just the final task success (Fig.
-7).
-In our evaluation, each sub-goal is checked independently, meaning we record whether the policy managed to achieve each individual step.
-To this end, we found cases where the RGB and RGB-D BC policies attempted to perform the final motion of the task (such as pouring or cutting) before actually grasping the object, indicating a confusion in the sequence of actions.
+The rubric captures success on several intermediate steps, not just final task completion (Fig.
+8).
+In our evaluation, each sub-goal is checked independently, meaning we record whether the policy achieves each individual step.
+We found cases where the RGB and RGB-D BC policies attempted the final motion of the task (such as pouring or cutting) before grasping the object, indicating confusion in action sequencing.
 In contrast, our keypoint-based policy typically reaches the final sub-goal, with occasional last-step slippage (object slipping from the gripper), suggesting strong early-stage competence.
@@
 In particular, we downsampled the demonstrations to 50% and 25% of the full set and measured the success rates (Fig.
-8.
+9).
 The results show that our keypoint-based method remains highly effective even with far fewer demonstrations, whereas the image-based baselines degrade significantly as the training data is reduced.
@@
 Standard deviation bars show variability across tasks.
-IV-D Keypoints Selections Baselines
-After demonstrating that a 2D keypoint-based policy can be highly effective, we ask whether our affordance-driven keypoint selection is essential.
-We compare our approach with several alternative keypoint selection strategies, each using $K=19$ (15 on object and 4 on end-effector) keypoints and the same downstream policy architecture for fairness:
-YOLO [37] + DINO [29]: Use a YOLO object detector to generate bounding boxes, then sample DINO keypoint correspondences within each detected box.
-YOLO [37] + Uniform Grid: Uniformly sample a grid of points inside each YOLO-detected bounding box.
-DINO [29] Only: Use DINO to generate correspondences over the entire image without bounding-box or filtering.
-SAM [38] + DINO: Use the Segment Anything Model (SAM) to obtain segmentation masks, use YOLO to identify the target object’s mask, and then sample DINO correspondences within that segmented region.
-Hand-Picked (P3PO): An oracle baseline where an expert manually chooses the keypoints (following P3PO [6]), and these points are then tracked using DIFT + CoTracker.
-Fig.
-9 summarizes the success rates of these strategies, averaged across all tasks, for both seen and unseen object instances.
-We find that the methods relying on YOLO detection do not significantly improve generalization to unseen object categories.
-This is likely because they depend on known object class labels (e.g., always looking for a “cup”), which can be misguiding when the object geometry changes (for example, encountering a teapot that might not be recognized as a cup).
-It also can suffer in cluttered scenes when the boxes are not tight or when multiple objects are present.
-In contrast, our affordance-driven approach does not depend on specific object labels; it directly identifies task-relevant regions (interaction hotspots) via the affordance model.
-We also note that the manually selected keypoints (P3PO) yield a success rate nearly on par with Afford2Act.
-However, our method achieves this performance without the tedious manual effort of choosing points, and it offers the flexibility of using natural language to specify the task.
-IV-E Generalization Analysis
-Beyond evaluating overall success, we systematically test four axes of generalization to assess how well Afford2Act transfers to new situations:
+IV-E Generalization (E3): Analysis
+Beyond overall success rates, we evaluate transfer along four complementary axes to characterize how Afford2Act behaves under distribution shift:
   1. 1.
-    Unseen Object (Same Category): Test the policy on novel instances of the same category (e.g., a new mug for the “hold” task or a different brush for “brush_with”).
+    Unseen Instance (Same Category): Evaluate on novel instances from the training category (e.g., a different mug for hold or a different brush for brush_with).
   2. 2.
-    Unseen Category: Apply the learned skill to an object from a category not encountered during training (e.g., using a spatula to stir instead of a spoon).
+    Novel Category (Shared Functional Parts): Evaluate on objects from categories not seen during training that support the same interaction (e.g., stirring with a spatula instead of a spoon).
   3. 3.
-    Background Variation: Evaluate the policy in environments with different backgrounds or lighting conditions.
+    Scene Variation: Evaluate under changes in background and lighting.
   4. 4.
-    Distractors: Introduce additional objects or even moving agents (e.g., a person performing an unrelated action nearby) to test the policy’s ability to stay focused on the target task in the presence of distractions.
+    Distractors: Introduce additional objects and moving agents (e.g., a person performing an unrelated action nearby) to test robustness to task-irrelevant dynamics.
 Fig.
 10 illustrates these generalization scenarios and the corresponding performance.
-Our keypoint-based policy exhibits strong transferability to novel object instances within seen categories and even to objects from entirely new categories (essentially achieving zero-shot generalization).
+Our keypoint-based policy exhibits strong transfer to novel object instances within seen categories and also transfers to several new categories in our test scenarios, particularly when objects share similar functional parts, suggesting promising zero-shot generalization.
 It also remains robust under background changes and in cluttered environments.
-For example, in one challenging trial, the system successfully transferred the learned affordance to a teapot from a mug in a highly cluttered scene, correctly identifying the meaningful keypoints despite different geometry.
+For example, in a particularly challenging trial, a policy trained to lift a mug successfully lifted a teapot in a heavily cluttered scene, with the pipeline still localizing the correct interaction keypoints despite substantial changes in object geometry and background.
 Likewise, the policy proved resilient to dynamic distractions, even with a human actively moving nearby.
-Figure 9: Total Average success rate across keypoint selection methods.
-Figure 10: Top: Average success rate on four scenarios aggregated across all tasks.
-Bottom: Representative rollouts for each scenario.
+Figure 10: Representative rollouts for each scenario.
 The small inset labeled “Ref” shows the training reference object for that task.
+We report the success rate for each scenario across all experiments in parentheses.
 IV-F Ablation Studies
-We conducted additional ablations to highlight the contribution of different components of our pipeline:
+We conducted additional ablations to highlight the contributions of different components of our pipeline:
 Language Prompt Robustness: We varied the wording of the affordance prompt to ensure our approach is not overly sensitive to specific language.
 Qualitatively, we found that synonyms or related phrases produce similar affordance keypoints.
-For example, using “sweep” instead of “brush_with” or “pick_up” instead of “hold” resulted in nearly identical affordance predictions on the relevant objects, indicating the model robustly understands the intended action.
-Gating Network: We removed the gating network so that each keypoint token receives equal weighting.
+For example, using “brush_with” instead of “sweep” or “mix” instead of “stir” resulted in nearly identical affordance predictions on the relevant objects, indicating that the model robustly captures the intended action semantics.
+Gating Network: We removed the gating network so that each keypoint token receives equal weight.
 This modification significantly degraded performance (especially in cluttered scenes), reducing the overall success rate from $86\%$ to $52\%$.
-During trajectory rollout, the gating mechanism allows the policy to focus only on the keypoints necessary.
-Effective Keypoint Count: We analyzed the gating network’s behavior by computing the effective number of keypoints being used by the policy at each time step.
+During trajectory rollout, the gating mechanism allows the policy to focus only on the necessary keypoints.
+Effective Keypoint Count: We analyzed the gating network’s behavior by computing the effective number of keypoints used by the policy at each time step.
 Given the gating output weights $w_{i}$} for the $K$ keypoint tokens, we define $K_{\text{eff}}=1/\sum_{i=1}^{K}w_{i}^{2}$.
@@
 This implies that the gating network effectively attends to only a subset of the keypoints at any given time, which helps the policy concentrate on the most informative features for the task.
-Figure 11: Average effective number of keypoints used (over all time steps) for each task. This data was collected through offline inference on all 40 training demos for each task.
-IV-G Long-Horizon Skills via Two-Stage Re-Anchoring
-For more complex, multi-phase tasks, we integrate Large Model task planning stage into our framework.
-We first use a vision-language model (VLM) planner to detect when the first sub-goal of the task has been completed (e.g., when the robot has successfully grasped the object).
-Once this condition is detected, we re-anchor half of the keypoints by providing an updated affordance prompt for the next phase of the task.
-In practice, this means replacing roughly half of the original keypoints with new ones tailored to the subsequent sub-task.
-This re-anchoring allows the policy to adapt to the changed state of the scene and effectively execute the next stage in the trajectory.
-We experimented with this augmentation on the Stir task, achieving an impressive $90\%$ success rate, further showing promise with incorporating VLMs with affordance-based keypoint selection (see video).
+Figure 11: (a) Average effective number of keypoints used (over all time steps) for each task.
+This data was collected through offline inference on all 40 training demos for each task.
+(b) Failure case due to weak affordance masking, leading to the gripper colliding with the spatula.
+IV-G Two-Stage Re-Anchoring for Multi-Phase Tasks
+To examine whether affordance-grounded keypoints can extend beyond single-stage interactions, we conduct a preliminary study on the Stir task using two-stage re-anchoring.
+This task is explicitly multi-phase: the robot must first grasp the spoon, then move to the bowl, and finally execute the stirring motion.
+A vision-language model (VLM) detects completion of the initial grasping phase; after this trigger, we update the affordance prompt and re-anchor a subset of keypoints for the second phase.
+In this setting, the approach achieves a $90\%$ success rate, indicating that online keypoint refresh can be useful for multi-phase manipulation.
+Future work will evaluate this mechanism across additional multi-phase tasks, alternative phase-transition criteria, and longer-horizon interactions.
+IV-H Failure Cases and Limitations
+A primary failure mode arises when the affordance prior is inaccurate or incomplete.
+Because keypoint distillation is conditioned on the predicted actionable region, an imperfect mask can omit the true contact area or include distractors, leading to unstable keypoints and incorrect control (Fig.
+11).
+A second limitation is partial observability from a single RGB view: heavy occlusions or viewpoint ambiguity can break affordance localization and tracking even when the behavior is feasible.
+These failures mainly reflect perception reliability rather than the keypoint-based policy design, and could be mitigated with confidence-aware filtering, lightweight re-anchoring, and occlusion/mask-noise augmentation.
 V Conclusions and Future Work
-We introduced Afford2Act, a three-stage imitation learning pipeline that leverages semantically meaningful keypoints derived from affordance text prompts.
-Through six real-world experiments, we demonstrated that Afford2Act is not only data-efficient and performant, but also generalizes effectively to unseen objects, categories, backgrounds, and distractors.
-While our work represents the first attempt at semantically meaningful keypoint-based imitation learning, the selected keypoints remain fixed during inference and cannot be adapted dynamically to evolving contexts.
+We present Afford2Act, an imitation learning framework that uses language-conditioned affordance cues to distill a compact set of semantically grounded 2D keypoints for manipulation from a single external RGB view.
+Across six real-robot tasks, Afford2Act demonstrates data-efficient learning and robust performance under real-world variation, including unseen instances, novel categories with shared functional parts, background changes, and distractors.
+This indicates that affordance-grounded keypoints can serve as an effective lightweight state representation for manipulation.
+A promising direction is to allow the keypoint representation to update online as the scene evolves.
 As discussed in Sec.
-IV-G, integrating task planning with keypoint resampling is a promising future direction, potentially enabling flexible long-horizon policy execution.
+IV-G, integrating lightweight task planning with confidence-aware keypoint resampling could improve adaptability during long-horizon execution, particularly under changing visibility or contact.